A wagon has a mass of 20 kg and is pulled at an angle of 16 degrees from the horizontal by a motor with a force of 100 N. Because of this force, the wagon is moving at a constant, horizontal speed of 5.0 m every two seconds. What is the power output of the motor pulling the wagon?
1
Expert's answer
2020-05-08T16:27:47-0400
"P=Fv\cos{16}=(100)(2.5)\cos{16}=240\ W"